• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name s.frullani

    Log entry time 02:55:34 on April 5,2005

    Entry number 143264

    keyword=VDC wire spectra

    In the first run taken in this shift we observe that the wire spectrum
    in VDC left is lower that the reference while is higher in the VDC right. Looking back to previous runs we see that this behaviour
    started with run 3211 - change of kinematics for the right arm but the left arm is unchanged. In run 3209, the first run after the access of the last night to reboot downstair the ROC5, both spectra are lower that the reference.
    We have not found any reference in the logbook to this problem.